OpenStudy (anonymous):

what are the steps to solve -5=s-3

OpenStudy (texaschic101):

-5 = s - 3 --- we want s on one side and everything else on the other side...so we will first add 3 to both sides. -5 + 3 = s - 3 + 3 -- simplify -2 = s --- see how the 3's cancelled each other out on the right side

OpenStudy (anonymous):

yea thank you

OpenStudy (texaschic101):

not a problem :) happy to help

OpenStudy (anonymous):

5/13t=-9